# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v1

@@ 5.2 Averaging Leads to Robustness
We explore the other core property, averaging operation, of FL and its correlation with transfer robustness. To change the degree of averaging in FL, we modify the number of clients selected to average at each round. We use the centralized model and source model with 30 clients’ data as substitute models to perform the transfer attack. Both models are early stopped at 90% as mentioned in Section 3.1. We train our federated target model with 100 users following Section 3.1. We plot the relation of T.Rate and the number of users in (a) to (d) of Figure 3. Both CNN and ResNet50 exhibit a decreasing trend as the number of clients used to average increases. We also provide statistical testing to validate the correlation in Appendix C. With the Spearman correlation coefficient, we report a significant negative correlation in all four experiments (all p-values are less than .001.001.001.001). To further validate this observation in more practical evaluations, we perform the attack in different training stages, with different architectures and different aggregation methods, as shown in (e), (f) of Figure 3 and Figure 6 in Appendix D, where similar trends can be observed. This evidence demonstrates that averaging in FedAvg improves the robustness of the federated model against transfer attacks.
@@ 5.3 Discussion
We list the results of our investigations as the below take-home messages and their implications:
@@
6 Theoretical Evidence
Background. We first introduce background and notation. We use $(X,Y)$ to denote a dataset, where $X\in\mathbb{R}^{n\times p}$ and $Y\in\mathbb{R}^{n}$. We use $(x,y)$ to denote a sample following Section 2. We consider the problem of federated learning optimization model:
$\min_{\theta}\{l(\theta)$ $\triangleq\sum_{k=1}^{N}p_{k}l_{k}(\theta)\},\quad\text{where}\quad l_{k}(\theta)=\frac{1}{n_{k}}\sum_{j=1}^{n_{k}}l(f(x;\theta),y)$ (6)
@@ 6 Theoretical Evidence
$N$ is the total number devices and $n_{k}$ is the number of samples possessed by device $k$. $p_{k}$ is the weight of device $k$ with the constraint that $\sum p_{k}=1$. Following DBLP:conf/iclr/LiHYWZ20 . we assume the algorithm performs a total of $T$ stochastic gradient descent (SGD) iterations and each round of local training possesses $E$ iterations. $\frac{T}{E}$ is thus the total communication times. At each communication, a maximal number of $K$ devices will participate in the process. We quantify the degree of non-iid using $\Gamma=l_{\star}-\sum_{k=1}^{N}p_{k}l_{k\star}$ where $l_{\star}$ and $l_{k\star}$ is the minimum value of $l(\theta)$ and $l_{k}(\theta)$ repectively. We use the relative increase of adversarial loss demontis2018adversarial to measure the transferability for easier derivation:
$R(x,y,\theta,\theta^{\prime})=\frac{\nabla_{x}l(f^{\prime}(x;\theta^{\prime}),y)^{T}\nabla_{x}l(f(x;\theta),y)}{\|\nabla_{x}l(f^{\prime}(x;\theta^{\prime}),y)\|_{2}\|\nabla_{x}l(f(x;\theta),y)\|_{2}}$ (7)
We provide detail derivation in Appendix F. Following assumptions from DBLP:conf/iclr/LiHYWZ20 , which we detail about in Appendix G, and additional assumptions for adversarial transferability, we provides a low bound for $R$.
Assumption 1: $\partial f(x;\theta)/\partial x=\theta^{T}\rho(x,\theta)$ where $\rho(x,\theta)$ is an arbitrary function in the forms of $\rho(x,\theta):\mathbb{R}^{1\times p}\times\mathbb{R}^{p\times 1}\rightarrow\mathbb{R}$.
There are many functions following our standards (e.g. $\ell_{2}$-norm regularized linear regression, logistic regression and softmax classifier). For the rest of our discussion, we define $\nabla l(f(x;\theta),y)=\partial f(x;\theta)/\partial x$
Lemma 6.1.
With Assumption 1, we have
With Lemma 6.1, $\theta$ can be directly measured by the cosine similarity of the parameters $\theta$ and $\theta^{\prime}$. Furthermore, as we need to offer a discussion regarding multiple aspects of the model, such as the loss, the parameters, and the data, we follow the previous convention DBLP:conf/iclr/LiHYWZ20 to focus on a narrower scope the model family:
Assumption 2: Model $f$ is in the form of $f(\theta)=\sum_{x\in X}(x\theta)^{2}$.
Notice that this is not a significant deviation from previous studies DBLP:conf/iclr/LiHYWZ20 that focus on $f(\theta)=\theta^{T}A\theta$ for detailed investigation of the parameter behaviors.
Assumption 3: The covariance of the samples we study is positive semidefinite, i.e., $X^{T}X\succcurlyeq\mathbf{I}$
The following theorem also depends on assumptions made by DBLP:conf/iclr/LiHYWZ20 which is detailed in Appendix G (Assumption 4-8)
@@
$\mathbb{E}[R(x,y,\theta_{\star},\theta^{\prime})]\geq\dfrac{2\mu(\gamma+T-1)\theta_{\star}^{T}\theta_{\star}}{4(B+C)\kappa+\mu^{2}\gamma\kappa\mathbb{E}\|\theta_{1}-\theta_{\star}\|^{2}}$ (9)
where $L,\mu,\sigma_{k},G$ are defined in the assumptions, $\kappa=\frac{L}{\mu}$, $\gamma=\max\{8\kappa,E\}$ , $B=\sum_{k=1}^{N}p_{k}^{2}\sigma_{k}^{2}+6L\Gamma+8(E-1)^{2}G^{2}$ and $C=\frac{4}{K}E^{2}G^{2}$. $\theta_{1}$ is the parameter after one step update of SGD. $\theta_{\star}$ is the optimal parameter for centralized model.
@@
With Theorem 6.2, we can derive the lower bound of the expectation of $R(x,y,\theta_{\star},\theta^{\prime\prime})$ by setting $E=1$ and $K=1$, where $\theta^{\prime\prime}$ represents the centralized source model.
$\mathbb{E}[R(x,y,\theta_{\star},\theta^{\prime})]\geq\dfrac{2\mu(\gamma+T-1)\theta_{\star}^{T}\theta_{\star}}{4(\sum_{k=1}^{N}p_{k}^{2}\sigma_{k}^{2}+6L\Gamma+4G^{2})\kappa+\mu^{2}\gamma\kappa\mathbb{E}\|\theta_{1}-\theta_{\star}\|^{2}}$ (10)
@@
The difference between the lower bound of FL (Lemma 6.1) and the centralized model (Corollary 6.3) lies in the denominator. With FL model, $B+C=\sum_{k=1}^{N}p_{k}^{2}\sigma_{k}^{2}+6L\Gamma+8(E-1)^{2}G^{2}+\frac{4}{K}E^{2}G^{2}$, while centralized gives a smaller $B+C=\sum_{k=1}^{N}p_{k}^{2}\sigma_{k}^{2}+6L\Gamma+4G^{2}$, which leads to a larger lower bound. Thus, the transferability of adversarial examples generated by the surrogate centralized model to attack the federated model is less than that when attacking a centralized model.
@@
$\mathbb{E}_{x^{\prime},{y}}\mathbb{E}_{\mathcal{T}}[\|y-f_{n}(x^{\prime})\|^{2}_{2}]=\mathbb{E}_{x^{\prime},{y}}[\|y-\bar{f}(x^{\prime})\|^{2}_{2}]+\frac{1}{n}\mathbb{E}_{x^{\prime},\mathcal{T}}[\|\bar{f}(x^{\prime})-f(x^{\prime},\mathcal{T})\|^{2}_{2}]$ (11)
@@ 7 Conclusion
This paper explores the potential for malicious clients to masquerade as benign entities in Federated Learning, then exploit this position to launch transferable attacks. Our evaluation shows that limited data can yield a comparable transfer rate to a full-dataset attack. We discover that decentralized training, heterogeneous data, and averaging operations enhance transfer robustness and reduce the transferability of adversarial examples.
Ethics Statement
In this work, we aim to further understand the security challenges a Federated Learning system can face in reality with a new and practical settings. We believe our study will benefit the society with a full understanding of the working mechanism of the potential attacks in the scenario raised in the paper, and thus build the ground for continued study of the methods that are secure against such threats.
Appendix A Comparison with query-based attack
Query-based black-box attack aims to attack the target model with limited queries to the target model. Through these queries, the source model manage to optimize the decision boundary towards the target model. However, most APIs and FL-based systems requires charges to access the service or equips anomaly detection that detects multiple or malicious queries. In this section, we provide a comparison of our proposed attack setting and the query-based black-box attack. To conduct a fair comparison, we follow the experiment setting in Section 4.2. With data from a limited number of users, instead of using the ground-truth class labels to train the source model, we query the target model for the prediction and set the predicted label as ground-truth. Note that, to facilitate the queries for the data from one client, one will have to perform 500 (in 100 clients partition) requests to the target model. The comparison is shown in Figure 4. We can see from Figure 4 that query-based outperforms the our proposed attack setting by a slight margin when the number of clients used to train the source model is small and the transfer rate of query-based attack continues to increase after 20 users where our attack setting begins to decrease. This can be attributed to the queries which help the source model to learn a more similar decision boundary. Through this experiment, we can hypothesize that, with some limited number of queries, our attack setting can be further boosted since it can help the substitute model to learn a similar loss landscape and decision boundary as the ones of target model.
@@ Appendix A Comparison with query-based attack
Figure 4: Comparison with query-based black-box attack; Left: transfer accuracy as a function of number of users’ data leveraged in source model training; Right: transfer rate as a function of number of users’ data leveraged in source model training;
Appendix B Heterogeneity in pathological non-iid distribution
To prove generalization of our findings with other non-iid distribution, we follow [30] to generate pathological distribution (partition by shards). To partition CIFAR10 datatset into 100 clients, we first sort the samples by class label and then divide the data into $n$ shards of size $50000/n$ (e.g. 500 shards of size 100) and assign each client $n/100$ shards. In this way, most client will have only samples of $n/100$ classes [30]. By varying the number of shards $n$, we control the maximum number of classes in most clients and hence the degree of non-iid of partitions. We conduct experiments on two different settings, the 10 user partition and 100 user partition. For 10 user partition, we generate a partition comprised of 10 users and vary the number of shards $n$ from 20 to 70. For 100 user partition, we generate a partition comprised of 10 users and vary the number of shards $n$ from 200 to 700. We generate the transfer rate v.s. maximum number of classes plots as in the Section 5.1. We can observe from Figure 5 a increase trend in both experiment settings with ResNet50 and CNN which demonstrates that our findings hold under different ways of simulating data heterogeneity. We also perform hypothesis testing for this correlation in Appendix C. All four experiments show significant correlation under a level of 0.01.
Figure 5: Transfer rate v.s. maximum number of classes per client; (a): Results of ResNet50 on CIFAR10 dataset; (b): Results of CNN on CIFAR10 dataset;
Notice we also find that the transfer rate of 10 users experiment is consistently larger than the 100 users setting which further demonstrates our findings in Section 5.1 that more decentralized training leads to lower adversarial transferability for federated model.
@@ Appendix C Statistical Hypothesis Testing on Spearman correlation coefficient
We demonstrates the correlation between different degree of decentralization, heterogeneity of training data and transferability of adversarial examples in Section 5.1 through plots and graphs and show that with more decentralized, heterogeneous data, the federated model is more robust to transfer attack. We also display a negative relation between the number of clients to average and transfer success rate through box charts in Section 5.2.
To statistically validate these correlations, we perform two-tailed Hypothesis Testing on Spearman correlation coefficient. To conduct Hypothesis Testing for Spearman correlation coefficient on a specified correlation, we first calculate the Spearman correlation coefficient $\rho$ on the two sets of points (e.g. T.Rate and Dirichlet $\alpha$):
@@
Appendix D Different aggregation method
To further validate the observation in Section 5.2 and see if this phenomenon generalizes to different aggregation methods. We perform the same experiment on three different aggregation methods, namely Krum [6], Geometric Mean [44] and Trimmed Mean [44]. As per Figure 6, we can see that with all three aggregation methods, there are decreasing trends as the number of averaged users per round increases similar as Figure 3. This further demonstrates our observations in Section 5.2 and support theoretical results in Section 6.
Figure 6: Transfer rate v.s. different number of clients selected to average in each round of difference aggregation method.
Appendix E Does different partition matters in transfer attack?
In Section 4.2, we discover that with source model trained in federated manner, T.Rate can be boosted to the highest of 99%. Since we assume that the attacker have information regarding the data partition of all malicious clients, we leverage the same partition used in target model to train the source model. In this section, we are curious about whether different partitions used by source model effects the transferability of its adversarial examples against the target model. That is, whether using a partition different from the target model to train the source model affects the transfer rate of its adversarial examples. To explore such setting, we first randomly generates two different partitions with distinct random seeds and then perform the source model training and transfer attack the same as in Section 4.2. We repeat the experiment 4 times with different random seeds and report the mean results in Figure 7. We observe no significant difference between the same partition and the different partition setting. To further validate this observation, we perform Hypothesis Test on the obtained results with Paired Sample T-Test. The p-value .393.393.393.393 means that there is no significant difference between the transfer rate obtained by same and different partition. This investigation and result further improve the possibility of attacking a FL system through black-box attack.
Figure 7: Difference between same partition and different partition; Left: transfer accuracy v.s. number of users’ data leveraged in source model training; Right: transfer rate v.s. number of users’ data leveraged in source model training;
Appendix F Relative Increase R
For simplicity, we derive another metric for theoretical analysis. We follow [11] and define the transferability to be the loss of the target model of an adversarial example, which can be simplified through a linear approximation of the loss function:
$T$ $=l(f(x+\hat{\delta};\theta),y)$ (12)
$\approx l(f(x;\theta),y)+\hat{\delta}^{T}\nabla_{x}l(f(x;\theta),y)$ (13)
where $\hat{\delta}$ is some perturbation generated by the surrogate model, which corresponds to the maximization of an inner product over an $\epsilon$-sized ball under the above linear approximation:
$\hat{\delta}$ $\in\operatorname*{arg\,max}_{\|\delta\|_{p}<\epsilon}l(f^{\prime}(x+\delta;\theta^{\prime}),y)$ (14)
$\max_{\|\delta\|_{p}<\epsilon}$ $\delta^{T}\nabla_{x}l(f^{\prime}(x;\theta^{\prime}),y)=\epsilon\|\nabla_{x}l(f^{\prime}(x;\theta^{\prime}),y)\|_{q}$ (15)
where $\|\|_{q}$ and $\|\|_{p}$ are dual norm.
Without loss of generality, we take $p=2$ and gives optimal $\hat{\delta}=\epsilon\frac{\nabla_{x}l(f^{\prime}(x;\theta^{\prime}),y)}{\|\nabla_{x}l(f^{\prime}(x;\theta^{\prime}),y)\|_{2}}$ from the surrogate model. Substituting it back to the Equation 13 we have the loss increment, bounded by the loss of the white-box attack:
$\Delta l=\epsilon\frac{\nabla_{x}l(f^{\prime}(x;\theta^{\prime})^{T}\nabla_{x}l(f(x;\theta),y)}{\|\nabla_{x}l(f^{\prime}(x;\theta^{\prime}),y)^{T}\|_{2}}\leq\epsilon\|\nabla_{x}l(f(x;\theta),y)\|_{2}$ (16)
We define the relative increase in loss in black-box case compared to white box as $R(x,y)$ which we show has a lower bound.
$R(x,y,\theta,\theta^{\prime})=\frac{\nabla_{x}l(f^{\prime}(x;\theta^{\prime}),y)^{T}\nabla_{x}l(f(x;\theta),y)}{\|\nabla_{x}l(f^{\prime}(x;\theta^{\prime}),y)\|_{2}\|\nabla_{x}l(f(x;\theta),y)\|_{2}}$ (17)
Appendix G Proof of Theorem 6.2
We give the rest of assumptions for Theorem 6.2 which is taken from [26].
@@ Appendix F Relative Increase R
We use Theorem 2 from [26] as our lemma to prove our Theorem 6.2.
Lemma G.1.
@@
Appendix H Proof of Theorem 6.6
@@
We adopt the Bias-Variance decomposition [16] to provide theoretical justification for the observation in Section 5.2. Let $\mathcal{T}$ denote the train set $\mathcal{T}=(X,Y)$ sampled from some data distribution and $x^{\prime}$ denote the adversarial example following some adversarial distribution. First, we give the Bias-Variance decomposition as follows:
$\mathbb{E}_{x^{\prime},{y}}\mathbb{E}_{\mathcal{T}}[\|y-f(x^{\prime};\mathcal{T})\|^{2}_{2}]$ $=\underbrace{\mathbb{E}_{x^{\prime},{y}}[\|y-\bar{f}(x^{\prime})\|^{2}_{2}]}_{\text{Bias}^{2}}+\underbrace{\mathbb{E}_{x^{\prime},\mathcal{T}}[\|\bar{f}(x^{\prime})-f(x^{\prime};\mathcal{T})\|^{2}_{2}]}_{\text{Variance}}$ (29)

### v2

@@ 5.2 Averaging Leads to Robustness
We explore the other core property, averaging operation, of FL and its correlation with transfer robustness. To change the degree of averaging in FL, we alter the number of clients selected to average at each round. To comprehensively illustrate this finding, we use two source models to attack. The first one is a centralized model trained with all the data. The Second surrogate model is trained with 30% of all the clients’ data, simulating the attack in Sec. 4.3. We plot the relation of T.Rate and #averaged users in (a) to (d) of Fig. 6. Both CNN and ResNet50 exhibit a decreasing trend as more users participate in the averaging operation. This demonstrates that the averaging operation contributes to the robustness of the FL model and more clients to average per round leads to higher transfer robustness. We provide statistical testing to validate the correlation in Appendix C. With the Spearman correlation coefficient, we report a significant negative correlation in all four experiments (all p-values are less than .001.001.001.001).
To provide a more practical evaluation, we follow the setting in practical evaluation of Sec. 4.3 and evaluate the above findings in different training stages and with different architectures. As shown in (e), (f) of Fig. 6, we can observe a similar trend as mentioned above. This demonstrates that this phenomena hold to more practical settings and wider configurations. We further illustrate that this observation generalizes to different aggregation methods in Appendix B.
Figure 5: Additional experiments on SVHN, CIFAR100 and ImageNet200. (a) Our attack with ResNet18 on SVHN (first row), ResNet50 on CIFAR100 (second row) and ResNet50 on ImageNet200 (third row). (b) How decentralization relates to the transfer robustness of FL model on SVHN (first row), CIFAR100 (second row) and ImageNet200 (third row) datasets. (c) How heterogeneity relates to the transfer robustness of FL model on SVHN (first row), CIFAR100 (second row) and ImageNet200 (third row) datasets. (d) How averaging operation relates to the transfer robustness of FL model on SVHN dataset (first row), CIFAR100 (second row) and ImageNet200 (third row) datasets.
@@ 5.3 Discussion
We summarize the above investigations as the below take-home messages and provide implications for understanding adversarial transferability in FL and secure FL applications:
@@
6 Supporting Theoretical Evidence
Notation.
@@ 6 Supporting Theoretical Evidence
We use $(X,Y)$ to denote a dataset, where $X\in\mathbb{R}^{n\times p}$ and $Y\in\mathbb{R}^{n}$. We use $(x,y)$ to denote a sample following Section 2. We consider the problem of federated learning optimization model: $\min_{\theta}\{l(\theta)\triangleq\sum_{k=1}^{N}p_{k}l_{k}(\theta)\},\quad$ where $\quad l_{k}(\theta)=\frac{1}{n_{k}}\sum_{j=1}^{n_{k}}l(f(x;\theta),y)$, $N$ is the total number devices and $n_{k}$ is the number of samples possessed by device $k$. $p_{k}$ is the weight of device $k$ with the constraint that $\sum p_{k}=1$. Following Li et al. (2020b). we assume the algorithm performs a total of $T$ stochastic gradient descent (SGD) iterations and each round of local training possesses $E$ iterations. $\frac{T}{E}$ is thus the total communication times. At each communication, a maximal number of $K$ devices will participate in the process. We quantify the degree of non-iid using $\Gamma=l_{\star}-\sum_{k=1}^{N}p_{k}l_{k\star}$ where $l_{\star}$ and $l_{k\star}$ is the minimum value of $l(\theta)$ and $l_{k}(\theta)$ respectively. We use the relative increase of adversarial loss Demontis et al. (2018) to measure the transferability for easier derivation, which is defined as the loss of the target model of an adversarial example, simplified through a linear approximation of the loss function: $\small T=l(f(x+\hat{\delta};\theta),y)\small\approx l(f(x;\theta),y)+\hat{\delta}^{T}\nabla_{x}l(f(x;\theta),y)$, where $\hat{\delta}$ is some perturbation generated by the surrogate model, which corresponds to the maximization of an inner product over an $\epsilon$-sized ball under the above linear approximation:
$\small\hat{\delta}\in\operatorname*{arg\,max}_{\|\delta\|_{p}<\epsilon}l(f^{\prime}(x+\delta;\theta^{\prime}),y),\quad\small\max_{\|\delta\|_{p}<\epsilon}\delta^{T}\nabla_{x}l(f^{\prime}(x;\theta^{\prime}),y)=\epsilon\|\nabla_{x}l(f^{\prime}(x;\theta^{\prime}),y)\|_{q}$
where $\|\|_{q}$ and $\|\|_{p}$ are dual norm.
Without loss of generality, we take $p=2$ and gives optimal $\hat{\delta}=\epsilon\nabla_{x}l(f^{\prime}(x;\theta^{\prime}),y)/\|\nabla_{x}l(f^{\prime}(x;\theta^{\prime}),y)\|_{2}$ from the surrogate model. Substituting it back to the Equation 6 we have the loss increment, bounded by the loss of the white-box attack:
$\small\Delta l=\epsilon\frac{\nabla_{x}l(f^{\prime}(x;\theta^{\prime})^{T}\nabla_{x}l(f(x;\theta),y)}{\|\nabla_{x}l(f^{\prime}(x;\theta^{\prime}),y)^{T}\|_{2}}\leq\epsilon\|\nabla_{x}l(f(x;\theta),y)\|_{2}$
We define the relative increase in loss in the black-box case compared to the white box as $R(x,y)$, which we show has a lower bound.
$\small R(x,y,\theta,\theta^{\prime})=\frac{\nabla_{x}l(f^{\prime}(x;\theta^{\prime}),y)^{T}\nabla_{x}l(f(x;\theta),y)}{\|\nabla_{x}l(f^{\prime}(x;\theta^{\prime}),y)\|_{2}\|\nabla_{x}l(f(x;\theta),y)\|_{2}}$ (5)
Then we provide a low bound for $R$.
@@
$\small\mathbb{E}[R(x,y,\theta_{\star},\theta^{\prime})]\geq\dfrac{2\mu(\gamma+T-1)\theta_{\star}^{T}\theta_{\star}}{4(B+C)\kappa+\mu^{2}\gamma\kappa\mathbb{E}\|\theta_{1}-\theta_{\star}\|^{2}}$ (6)
where $L,\mu,\sigma_{k},G$ are defined in the assumptions, $\kappa=\frac{L}{\mu}$, $\gamma=\max\{8\kappa,E\}$ , $B=\sum_{k=1}^{N}p_{k}^{2}\sigma_{k}^{2}+6L\Gamma+8(E-1)^{2}G^{2}$ and $C=\frac{4}{K}E^{2}G^{2}$. $\theta_{1}$ is the parameter after one step update of SGD. $\theta_{\star}$ is the optimal parameter for the centralized model.
We refer to Appendix D for proof.
@@
We derive the lower bound of the expectation of $R(x,y,\theta_{\star},\theta^{\prime\prime})$ by setting $E=1$ and $K=1$, where $\theta^{\prime\prime}$ represents the centralized source model.
$\mathbb{E}[R(x,y,\theta_{\star},\theta^{\prime})]\geq\dfrac{2\mu(\gamma+T-1)\theta_{\star}^{T}\theta_{\star}}{4(\sum_{k=1}^{N}p_{k}^{2}\sigma_{k}^{2}+6L\Gamma+4G^{2})\kappa+\mu^{2}\gamma\kappa\mathbb{E}\|\theta_{1}-\theta_{\star}\|^{2}}$
@@
The difference between the lower bound of FL (Lemma D.1) and the centralized model (Corollary 6.2) lies in the denominator. With FL model,
$\small B+C=\sum_{k=1}^{N}p_{k}^{2}\sigma_{k}^{2}+6L\Gamma+8(E-1)^{2}G^{2}+\frac{4}{K}E^{2}G^{2},$
while centralized gives a smaller
$\small B+C=\sum_{k=1}^{N}p_{k}^{2}\sigma_{k}^{2}+6L\Gamma+4G^{2},$
leading to a larger lower bound. Thus, the transferability of adversarial examples generated by the surrogate centralized model to attack the federated model is less than that when attacking a centralized model.
@@
$\mathbb{E}_{x^{\prime},{y}}\mathbb{E}_{\mathcal{T}}[\|y-f_{n}(x^{\prime})\|^{2}_{2}]=\mathbb{E}_{x^{\prime},{y}}[\|y-\bar{f}(x^{\prime})\|^{2}_{2}]+\frac{1}{n}\mathbb{E}_{x^{\prime},\mathcal{T}}[\|\bar{f}(x^{\prime})-f(x^{\prime},\mathcal{T})\|^{2}_{2}]$
@@ 7 Conclusion
We explore the potential for malicious clients to masquerade as benign entities in Federated Learning, then exploit this position to launch transferable attacks. We provide a thorough investigation of the proposed attack with limited data setting. Our evaluation shows that limited data can yield a comparable transfer rate to a full-dataset attack.
To fully understand how adversarial examples transfer between centralized and federated models, we further study two intrinsic properties of FL and its relation with transfer robustness. We discover that decentralized training, heterogeneous data, and averaging operations enhance transfer robustness and reduce the transferability of adversarial examples. We provide evidence from both the perspective of empirical experiments and theoretical analysis.
Our findings have implications for understanding the robustness of federated learning systems and poses a practical question for federated learning applications.
Appendix A Attack with Same or Different Partition
We show in Section 4.3 that training surrogate models in a federated manner can lead to even higher transferability of adversarial examples. However, in the experiment in Section 4.3, we partition the collected data from malicious clients as the federated model and train the source model in a federated manner. In this section, we further evaluate whether this partition information is crucial to the higher transferability. That is, whether different partitions used by the source model affect the transferability of its adversarial examples against the target model.
@@ Appendix A Attack with Same or Different Partition
Figure 7: Difference between same partition and different partition; Left: transfer accuracy v.s. number of users’ data leveraged in source model training; Right: transfer rate v.s. number of users’ data leveraged in source model training
To simulate this setting, we first randomly generate two different partitions with distinct random seeds and then perform the source model training and the target model training on the two different partitions. Then transfer attack is performed with the source model against the target model following the above configuration. We repeat the experiment 4 times with different random seeds and report the averaged results in Figure 7. We observe no significant difference between the same partition and the different partition settings. To further validate this observation, we perform a Hypothesis Test on the obtained results with the Paired Sample T-Test and achieve a p-value of .393.393.393.393 meaning that there is no significant difference. This further demonstrates the possibility of attacking a federated learning system through our proposed attacks and illustrates a higher security risk.
Appendix B Averaging Leads to Adversarial Robustness with Different Aggregation Methods
We show in Section 5.2 that the averaging operation contributes to the robustness of the FL model and more clients to average per round leads to higher transfer robustness. We further validate that this observation generalizes to different aggregation methods, i.e. Krum Blanchard et al. (2017), Geometric Mean Yin et al. (2018) and Trimmed Mean Yin et al. (2018). As per Figure 8, we can see that with all three aggregation methods, there are decreasing trends as the number of averaged users per round increases.
Figure 8: Transfer rate v.s. #clients selected to average in each round with difference aggregation methods.
@@ Appendix C Statistical Hypothesis Testing on Spearman correlation coefficient
In Section 5.1, we elucidate the correlation between the degree of decentralization and the heterogeneity of training data with the transferability of adversarial examples. The findings denote that, with more decentralized, heterogeneous data, the federated model is more robust to transfer attack. Furthermore, Section 5.2 portrays a discernible inverse relationship between the number of clients and the average transfer success rate, as depicted through box plots, which illustrate the averaging operation leads to better robustness against adversarial examples.
To statistically validate these correlations, we perform two-tailed Hypothesis Testing on Spearman correlation coefficient. To conduct Hypothesis Testing for Spearman correlation coefficient on a specified correlation, we first calculate the Spearman correlation coefficient $\rho$ on the two sets of points (e.g., T.Rate and Dirichlet $\alpha$):
@@
Appendix D Proof of Theorem 6.1
Following assumptions from Li et al. (2020b) and additional assumptions for adversarial transferability, we provide a low bound for $R$.
Assumption 1
$\partial f(x;\theta)/\partial x=\theta^{T}\rho(x,\theta)$ where $\rho(x,\theta)$ is an arbitrary function in the forms of $\rho(x,\theta):\mathbb{R}^{1\times p}\times\mathbb{R}^{p\times 1}\rightarrow\mathbb{R}$.
There are many functions following our standards (e.g., $\ell_{2}$-norm regularized linear regression, logistic regression and softmax classifier). For the rest of our discussion, we define $\nabla l(f(x;\theta),y)=\partial f(x;\theta)/\partial x$
Lemma D.1.
With Assumption 1, we have
With Lemma D.1, $\theta$ can be directly measured by the cosine similarity of the parameters $\theta$ and $\theta^{\prime}$. Furthermore, as we need to offer a discussion regarding multiple aspects of the model, such as the loss, the parameters, and the data, we follow the previous convention Li et al. (2020b) to focus on a narrower scope the model family:
Assumption 2
Model $f$ is in the form of $f(\theta)=\sum_{x\in X}(x\theta)^{2}$
Notice that this is not a significant deviation from previous studies Li et al. (2020b) that focus on $f(\theta)=\theta^{T}A\theta$ for detailed investigation of the parameter behaviors.
Assumption 3
The covariance of the samples we study is positive semidefinite, i.e., $X^{T}X\succcurlyeq\mathbf{I}$
@@
We use Theorem 2 from Li et al. (2020b) as our lemma to prove our Theorem 6.1.
Lemma D.2.
@@
Appendix E Proof of Theorem 6.5
@@
We adopt the Bias-Variance decomposition Geman et al. (1992) to provide theoretical justification for the observation in Section 5.2. Let $\mathcal{T}$ denote the train set $\mathcal{T}=(X,Y)$ sampled from some data distribution and $x^{\prime}$ denote the adversarial example following some adversarial distribution. First, we give the Bias-Variance decomposition as follows:
$\mathbb{E}_{x^{\prime},{y}}\mathbb{E}_{\mathcal{T}}[\|y-f(x^{\prime};\mathcal{T})\|^{2}_{2}]$ $=\underbrace{\mathbb{E}_{x^{\prime},{y}}[\|y-\bar{f}(x^{\prime})\|^{2}_{2}]}_{\text{Bias}^{2}}$ (19)
$+\underbrace{\mathbb{E}_{x^{\prime},\mathcal{T}}[\|\bar{f}(x^{\prime})-f(x^{\prime};\mathcal{T})\|^{2}_{2}]}_{\text{Variance}}$